OpenStudy (anonymous):

so this is under solving inequalities,,.. this is the question, it is grade nine math chpter nine. a ---- > or equal to 3 -1.5

OpenStudy (anonymous):



OpenStudy (anonymous):

how did you find that????

OpenStudy (anonymous):

multiply -1.5 on both sides and you flip the sign

OpenStudy (unklerhaukus):

multiplied both sides -1.5, (the inequality symbol switches because we are multiplying by a negative)

OpenStudy (anonymous):

can you guys show me the work? like using a srawing?

OpenStudy (anonymous):

drawing*

OpenStudy (anonymous):

help!?

OpenStudy (saifoo.khan):

\[-a \ge3*1.5\] \[a \le -4.5\]

OpenStudy (akshay_budhkar):

\[{a \over -1.5} \ge 3\] multiplying by 1.5 on both the sides\[-a \ge 3 \times 1.5\] \[-a \ge 4.5\] now as it is a minus sign, we multiply by -1, when we do that the inequality sign changes so we have\[a \le -4.5\]

OpenStudy (akshay_budhkar):

is it clear now?

OpenStudy (anonymous):

yessss! ahh thanks :)

OpenStudy (akshay_budhkar):

good that you get it! my pleasure :D

OpenStudy (anonymous):

hold on, how would i show that the inequality holds true?

OpenStudy (anonymous):

hellllo???

OpenStudy (anonymous):

help!!!!!

OpenStudy (anonymous):

moneybird